In a certain reaction, 3.45g of Sodium reacted with 5.43g of Chlorine. What is the mass of the product formed in mg

A. 3.45 x 10-3 mg

B. 5.43 x 10-4 mg

C. 5.345 x 10-5 mg

D. 3.4 x 10-2 mg


1
Expert's answer
2021-11-08T08:11:31-0500

Na+2Cl-->>NaCl2

Molar mass of Na= 22.989769

3.45/22.989769= 0.15

Molar mass of Cl= 35.453

5.43/35.453= 0.1532

0.1532/2= 0.0766

= 0.0766+0.1532=0.2298

Molar mass of NaCl2= 58.44

= 0.2298×58.44

= . 3.4 x 10-2mg